• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 06:57:25 on October 4,2010

    Entry number 333587

    keyword=Owl shift summary

    Shift started after Moller commissioning.

    We checked that Compton rates were still too high to operate the detector.

    We proceeded to calibrate BPMs. The calibration was aborted before the end due to a failure in the harp scan software. The data taken is probably enough, but if we have some time we can continue some other day.

    We made 2 raster runs to calibrate its size.

    We took a very short run on the multifoil target with 2uA to test backgrounds on the calorimeter. However, the run was stopped early because MCC wanted to send tune beam to reset some wrong setting on some Compton magnets.

    At 6:45 we turned the beam over to MCC for them to perform the multiharp scan tests.

    At 7:00, a BCM calibration is planned. The new crew will take care of it.